# Neuro-Informed Joint Learning Enhances Cognitive Workload Decoding in Portable BCIs

### v1

@@ III-B Analysis and Discussion
The comparative analysis reveals distinct performance advantages of our proposed MuseCogNet. As seen in Table I, established models like EEGNet [8] and DeepConvNet [9] demonstrate competent baseline performance (Accuracy: 57.07%, F1: 51.25-52.09%), validating their status as canonical architectures for EEG decoding. Meanwhile, our implementation of ResNet achieves 58.41% accuracy, closely matching the original CLDrive benchmark(58.13%) [7], confirming proper experimental replication. Critically, MuseCogNet achieves superior performance of 62.68% accuracy and 57.24% F1, representing absolute gains of 4.27% accuracy and 5.63% F1 over the strongest baseline (ResNet). We attribute this significant margin to our novel joint learning paradigm enhanced by average-pooled reconstruction objective, where the SSL component learns steady neurodynamic patterns that conventional supervised CNNs fail to capture. The ablation study (Table II) provides definitive evidence: Removing SSL components causes performance degradation of 1.91% accuracy and 2.64% F1, confirming that our average-pooled reconstruction objective contributes substantially to feature robustness.

### v2

@@ III-B Analysis and Discussion
The comparative analysis reveals distinct performance advantages of our proposed MuseCogNet. As seen in Table I, established models like EEGNet [8] and Deep ConvNet [9] demonstrate competent baseline performance (Accuracy: 57.07%, F1: 51.25-52.09%), validating their status as canonical architectures for EEG decoding. Meanwhile, our implementation of ResNet achieves 58.41% accuracy, closely matching the original CL-Drive benchmark(58.13%) [7], confirming proper experimental replication. Critically, MuseCogNet achieves superior performance of 62.68% accuracy and 57.24% F1, representing absolute gains of 4.27% accuracy and 5.63% F1 over the strongest baseline (ResNet). We attribute this significant margin to our novel joint learning paradigm enhanced by average-pooled reconstruction objective, where the SSL component learns steady neurodynamic patterns that conventional supervised CNNs fail to capture. The ablation study (Table II) provides definitive evidence: Removing SSL components causes performance degradation of 1.91% accuracy and 2.64% F1, confirming that our average-pooled reconstruction objective contributes substantially to feature robustness.